OpenStudy (anonymous):

Find the function rule X: -2, -1,0, 1, 2 Y: 9, 4 -1,-6,-11

OpenStudy (lucaz):

notice as x increases by 1, y decreases by 5, the slope is -5, and when x=0, y=-1 so in the slope-intercept form y=mx+b, where m is the slope and b is the value of y when x=0 you have: f(x) or y=-5(x)+(-1) y=-5x-1

OpenStudy (lucaz):

slope = (change in y) / (change in x)

OpenStudy (anonymous):

so the rule is?

OpenStudy (lucaz):

this equation y=-5x-1

OpenStudy (anonymous):

Thank you, i became a fan :)

OpenStudy (lucaz):

you're welcome